### v1

@@ 3.1.5 Diverse Strategies for Enhancing Domain Adaptation
The adaptation techniques can be further nuanced based on class overlap scenarios between the source and target domains, leading to Closed-Set, Partial-Set, and Open-Set Adaptation (CPO). Each addresses a specific kind of overlap, from complete to none, reflecting the diverse challenges in domain adaptation. Pseudo Label Refinement(PLR) , on the other hand, enhances the reliability of labels in unsupervised settings, refining the model-generated labels for better accuracy.
Lastly, Contrastive Learning (CL), by distinguishing between similar and dissimilar data points, offers a robust way for models to learn distinctive features, essential for tasks like classification and clustering. For methods like game theory, which are sparsely used, we have categorized them under Aditional Emerging Methods (AEM) to provide a comprehensive overview.
@@ 3.2 Vision Transformers in Domain Generalization
In our comprehensive review of the existing research, we analyzed how ViTs are adapted for the DG process. Based on our findings, we have identified four distinct approaches that encapsulate the common strategies within the literature. we categorized the approaches into four main categories based on our analysis of the literature. These categories are: Multi-Domain Learning, Meta-Learning Approaches, Regularization Techniques, and Data Augmentation Strategies. In the subsequent section, we will delve into the specifics of the research within each category.
@@ 3.2.2 Meta-Learning Approaches
Meta-learning is an approach for training ViTs to adapt rapidly to new domains with minimal data. By engaging in a variety of learning tasks, ViTs develop the ability to apply meta-knowledge across different settings, significantly boosting their adaptability and performance in unseen environments. In the following paragraphs, we will delve into the details of research focusing on meta-learning approaches.
[105] introduces the DoPrompt algorithm, a novel approach in the realm of ViTs for domain generalization. It uniquely incorporates Domain Prompt Learning (DPL) and Prompt Adapter Learning (PAL), embedding domain-specific knowledge into prompts for each source domain. These prompts are then integrated through a prompt adapter for effective target domain prediction.
@@
Researchers in [107], propose Common-Specific Visual Prompt Tuning (CSVPT), a new method integrating domain-common prompts to capture task context and sample-specific prompts to address data distribution variations, enabled by a trainable prompt-generating module (PGM). This approach is specifically tailored for effective adaptation to unknown testing domains, significantly enhancing out-of-distribution generalization in image classification tasks.
@@ 3.2.3 Regularization Techniques
Regularization methods are identified as essential for preventing overfitting and promoting the learning of generalized features.
@@ 3.2.4 Data Augmentation Strategies
Advanced data augmentation techniques, including synthetic data generation, are employed to broaden the training dataset’s diversity. This strategy is instrumental in preparing ViTs to better generalize to novel domains. Upcoming segments will illustrate different data augmentation strategies employed in ViTs.
@@
In conclusion, table 3 presents representative works from recent research that have employed ViTs for DG. These selected studies highlight the adaptability and potential of ViTs in improving models’ ability to generalize across various domains.
@@
Category
Method
Design Highlights
Training Strategies
Publication
Meta-Learning Approaches
CSVPT[107]
Boosts OOD generalization with dynamically generated domain-invariant and variant prompts via a trainable module, improving adaptability across datasets
Cross Entropy
ACCV 2022
Regularization Techniques
SDViT[108]
Reduces overfitting by using self-distillation, entropy-based signals, and a modular approach, aiming for better learning across different domains
Cross Entropy, KL Divergence
ACCV 2022
T3A[111]
Enhances DG by updating linear classifiers with online-generated pseudo-prototypes, offering robustness in varying environments without back-propagation
Cross Entropy, Pseudo Label Refinement
NeurIPS 2021
@@ 3.2.5 Diverse Strategies for Enhancing Domain Generalization
Table 4: Comprehensive Summary of Techniques in ViTs for DG: DST (Domain Synthesis), SD (Self Distillation), CGFL (Class Guided Feature Learning), ADPL (Adaptive Learning), ViT-CNN (Hybrid Networks), FAug (Feature Augmentation/Feature Learning), PL (Prompt-Learning), CRD( Cross Domain), ViT+CNN(Hybrid Networks), SDKD (Source Domain Knowledge Decding), KD (Knowledge Distillation), SFDA (Source-Free DA), MML (Multi Modal Learning), and TTA (Test-Time Adaptation)
@@ 4.2 Action Recognition
Lin et al. [138] introduce ViTTA, a method enhancing action recognition models during test time without retraining. This approach focuses on feature distribution alignment, dynamically adjusting to match test set statistics with those of the training set. A key aspect is its applicability to both convolutional and transformer-based networks. ViTTA also enforces consistency in predictions over temporally augmented video views, a strategy that significantly improves performance in scenarios with distribution shifts, showcasing its effectiveness over previous test-time adaptation techniques. Q. Yan and Y. Hu’s research [139] introduces a UDA method tailored for skeleton behavior recognition, addressing the challenge of aligning source and target datasets in domain adaptation. Their method employs a spatial-temporal transformer framework with three flows—source, target, and source-target facilitating effective domain alignment and handling variations in joint numbers and positions across datasets. Key to this approach is the use of subsequence encoding and an attention mechanism that emphasizes local joint relationships, thereby enhancing the representation of skeleton behavior. Comprehensive testing on various skeleton datasets shows the superiority of their Spatial-Temporal Transformer-based DA (STT-DA) method, underscoring its effectiveness in managing the complexities of domain adaptation in skeleton behavior recognition. The concept of applying transformers for skeleton-based action recognition in DA is viewed as a promising and potentially impactful direction in this field of study.[140]
@@ 4.3 Face Analysis
Face anti-spoofing (FAS) is a crucial aspect of biometric security systems, addressing the challenge of distinguishing between genuine and fake facial representations [141]. Recently WACV 2023 research [142]introduced a new approach for FAS using ViTs. The authors proposed the Domain-invariant ViT (DiVT) which employs two specific losses to enhance generalizability. The losses include a concentration loss for learning domain-invariant representations by aggregating features of real face data, and a separation loss to differentiate each type of attack across domains. The study highlights the effectiveness of transformers in capturing long-range dependencies and globally distributed cues, crucial for FAS tasks. It also addresses the large model size and computational resource issues commonly associated with transformer models by adopting a lightweight transformer model, MobileViT. The proposed approach differs from previous methods by focusing on the core characteristics of real faces’ features and unifying attack types across domains, leading to improved performance and efficiency in FAS applications. In addition, researchers in [141] explored FAS in surveillance contexts, where image quality varies widely. The paper introduces an Adversarial DG Network (ADGN), which classifies training data into sub-source domains based on image quality scores. It then employs adversarial learning for feature extraction and domain discrimination, achieving quality-invariant features. The approach also integrates transfer learning to mitigate limited training data issues. This innovative method proved effective in surveillance FAS, as evidenced by its performance in the 4th Face Anti-Spoofing Challenge at CVPR 2023.
@@ 4.4 Medical Imaging
Another challenging area is medical image segmentation, which struggles due to the inherent variability of medical images. To tackle the limited availability of training datasets, data-efficient ViTs were proposed. However, indiscriminate dataset combinations can result in Negative Knowledge Transfer (NKT). Addressing this, the introduction of MDViT, a multi-domain ViT with domain adapters, marked a significant advancement. This approach allowed for the effective utilization of varied data resources while mitigating NKT, showcasing superior segmentation performance even with an increase in domain diversity [148]. The robustness against adversarial attacks is a non-negotiable aspect of deep medical diagnosis systems. A novel CNN-Transformer hybrid model was introduced to bolster this robustness and enhance generalization. This model augmented image shape information in high-level feature spaces, smoothing decision boundaries and thereby improving performance on standardized datasets like MedMNIST-2D[149]. Liu et al.’s introduction of the Convolutional Swin-Unet (CS-Unet) represents a notable advance in medical image semantic segmentation. By integrating the Convolutional Swin Transformer (CST) block into transformers, they effectively combined multi-head self-attention with convolutions, providing localized spatial context and inductive biases essential for delineating organ boundaries. This model’s efficiency and effectiveness, particularly in its capability to surpass existing transformer-based methods without extensive pre-training, underscore the importance of localized spatial modeling in medical imaging [150]. A significant stride in domain generalization was achieved with the introduction of BigAug, a deep stacked transformation approach. This method applies extensive data augmentation, simulating domain shifts across MRI and ultrasound imaging. BigAug’s application of nine transformations to each image, validated across diverse segmentation tasks and challenge datasets, set a new standard, outperforming traditional augmentation and domain adaptation techniques in unseen domains [151].
@@
In reviewing the recent advancements in the application of ViTs for DA and Generalization in medical imaging, it becomes evident that ViTs are not only versatile but also increasingly effective in addressing domain-specific challenges. The studies surveyed indicate a significant shift towards models that are more adaptable to varying clinical environments, a crucial aspect of real-world medical applications. However, there is an observable need for further refinement in these models to ensure even greater accuracy and robustness, particularly in scenarios involving scarce or highly variable data. The success of models like EPVT and MDViT in enhancing generalization capabilities across diverse environments suggests a promising direction toward more domain-agnostic approaches. Nevertheless, the balance between model complexity and interpretability remains a key area for future exploration. As the field moves forward, there’s potential for integrating more advanced self-supervised learning techniques and exploring hybrid models that combine the strengths of both CNNs and ViTs. This could lead to a new generation of medical imaging tools that are not only more efficient in handling domain shifts but also more accessible and reliable for clinicians in varied healthcare settings.
@@
To conclude, this chapter has showcased the extensive utility of ViTs beyond image recognition tasks, highlighting their significant impact in areas such as semantic segmentation, action recognition, face analysis, and medical imaging. Beyond these fields, ViTs have proven to be highly adaptable and effective in sectors like precision agriculture and autonomous driving [155, 156]. These researches highlight that the potential of ViTs extends far beyond the initially discussed applications. Their adaptability to different environments and challenges showcases a growing research interest in diverse fields. Future research could explore even more innovative applications, leveraging ViTs’ unique ability to handle complex visual tasks and distribution shifts across different industries.
5 Conclusions and Future Research Directions
@@
This comprehensive review looks at how modern vision networks, ViTs, are used in DA and DG methods to handle distribution shifts. Reviewing the research, we’ve observed that the transformer’s self-attention mechanisms play a pivotal role in generalizing and adapting to new, distribution-shifted samples, as evidenced by reviewing the experiment results of research in the known shifted datasets including ImageNet-A, ImageNet-C, and Stylized ImageNet. With respect to distribution shifts’ handling methods, for DA methods, we’ve organized the research into four categories: feature-level, instance-level, model-level adaptations, and hybrid approaches. Additionally, we’ve introduced a new category for studies that combine different strategies with ViTs to tackle distribution shifts. These hybrid networks show the use of various strategies alongside ViTs, leading us to categorize the papers based on these combined approaches. A similar methodological framework was applied to DG, wherein papers are classified based on multi-domain learning, meta-learning, regularization techniques, and data augmentation strategies. These two distinct methods of categorization, by considering various approaches, provide deeper insights and a broader perspective. Finally, we extend our review to applications beyond image recognition, showcasing the versatility of these methods in various domains, and highlighting the potential of ViTs in a broad range of applications beyond the image recognition tasks.
@@
In our discussion and exploration of various research works, it becomes evident that we are at the nascent stage of developing these modern vision networks. Researchers are increasingly focusing on the characteristics of ViTs across diverse deep learning scenarios. While ViTs face challenges stemming from various factors, they are on a trajectory of improvement, much like CNNs were in their early stages. There is a noticeable rise in publications dedicated to modern vision models, emphasizing the advancements being made. As highlighted earlier, particularly due to the attention mechanism inherent in ViTs, their proficiency in handling distribution shifts is undeniable, marking an advantage over CNNs in this aspect. In line with the focus of this study, this section intends to outline potential research areas in the field of tackling distribution shifts. The recommendations take into account the properties of ViTs, as well as DA and DG approaches for managing distribution shifts, and how they can be effectively integrated. ViTs have demonstrated the capacity to surpass previous SOTA methods in certain contexts. However, their superiority is not consistently overwhelming, partly due to their reliance on manually designed architectures which may impede their adaptability, as explored by [50]

### v2

@@ 3.1.5 Diverse Strategies for Enhancing Domain Adaptation
The adaptation techniques can be further nuanced based on class overlap scenarios between the source and target domains, leading to Closed-Set, Partial-Set, and Open-Set Adaptation (CPO). Each addresses a specific kind of overlap, from complete to none, reflecting the diverse challenges in domain adaptation. Pseudo Label Refinement(PLR) , on the other hand, enhances the reliability of labels in unsupervised settings, refining the model-generated labels for better accuracy. Lastly, Contrastive Learning (CL), by distinguishing between similar and dissimilar data points, offers a robust way for models to learn distinctive features, essential for tasks like classification and clustering. For methods like game theory, which are sparsely used, we have categorized them under Aditional Emerging Methods (AEM) to provide a comprehensive overview.
@@ 3.2 Vision Transformers in Domain Generalization
In our comprehensive review of the existing researches, we analyzed how ViTs are adapted for the DG process. Based on our findings, we have identified four distinct approaches that encapsulate the common strategies within the literature. we categorized the approaches into four main categories based on our analysis of the literature. These categories are: Multi-Domain Learning, Meta-Learning Approaches, Regularization Techniques, and Data Augmentation Strategies. In the subsequent section, we will delve into the specifics of the research within each category.
@@ 3.2.2 Meta-Learning Approaches
Meta-learning is an approach for training ViTs to adapt rapidly to new domains with minimal data. By engaging in a variety of learning tasks, ViTs develop the ability to apply meta-knowledge across different settings, significantly boosting their adaptability and performance in unseen environments. We categorize several recent studies under meta-learning approaches, including Domain Prompt Learning (DPL) with the DoPrompt algorithm, hybrid architecture with query-memory decoding, and Common-Specific Visual Prompt Tuning (CSVPT), all of which illustrate the effectiveness of these techniques in improving domain generalization. In the following paragraphs, we will delve into the details of research focusing on meta-learning approaches.
[108] introduces the DoPrompt algorithm, a novel approach in the realm of ViTs for domain generalization. It uniquely incorporates Domain Prompt Learning and Prompt Adapter Learning, embedding domain-specific knowledge into prompts for each source domain. These prompts are then integrated through a prompt adapter for effective target domain prediction.
@@
Researchers in [110], propose Common-Specific Visual Prompt Tuning, a new method integrating domain-common prompts to capture task context and sample-specific prompts to address data distribution variations, enabled by a trainable prompt-generating module (PGM). This approach is specifically tailored for effective adaptation to unknown testing domains, significantly enhancing out-of-distribution generalization in image classification tasks.
@@ 3.2.3 Regularization Techniques
Regularization methods are essential for preventing overfitting and promoting the learning of generalized features. These methods impose various constraints during training to ensure that the learned features are not overly specific to the source domain, thus improving the model’s performance on unseen target domains. Regularization techniques such as self-distillation, cross-attention mechanisms, and test-time adjustments have been shown to significantly enhance the generalization capabilities of ViTs. By encouraging the model to learn broadly applicable features, these methods help to mitigate the impact of domain shifts.
@@ 3.2.4 Data Augmentation Strategies
Data augmentation strategies are applied to increase the diversity and robustness of training datasets. By artificially expanding the training data, these methods help ViTs to learn more generalized and adaptable features, improving their performance on unseen target domains. Advanced data augmentation techniques, including synthetic data generation, spatial transformation, and token-level feature stylization, have shown significant promise in enhancing the generalization capabilities of ViTs. By introducing variability in the training data, these methods help to mitigate the impact of domain shifts and improve model robustness.
@@
In conclusion, table LABEL:tab:Design_highlights_ViT_DG presents representative works from recent research that have employed ViTs for DG. These selected studies highlight the adaptability and potential of ViTs in improving models’ ability to generalize across various domains.
@@
Category Method Design Highlights Training Strategies Publication
Meta-Learning Approaches CSVPT[110] Boosts OOD generalization with dynamically generated domain-invariant and variant prompts via a trainable module, improving adaptability across datasets Cross Entropy ACCV 2022
Regularization Techniques SDViT[111] Reduces overfitting by using self-distillation, entropy-based signals, and a modular approach, aiming for better learning across different domains Cross Entropy, KL Divergence ACCV 2022
T3A[114] Enhances DG by updating linear classifiers with online-generated pseudo-prototypes, offering robustness in varying environments without back-propagation Cross Entropy, Pseudo Label Refinement NeurIPS 2021
@@ 3.2.5 Diverse Strategies for Enhancing Domain Generalization
To conclude this chapter, we have discussed various tables analyzing different ViT-based methods for DA and DG from multiple viewpoints. For a summary of the advantages and limitations of these methods, refer to Table 5 . The tables in this chapter offer a comprehensive overview and detailed analysis of the studies from different perspectives.
Table 4: Comprehensive Summary of Techniques in ViTs for DG: DST (Domain Synthesis), SD (Self Distillation), CGFL (Class Guided Feature Learning), ADPL (Adaptive Learning), ViT-CNN (Hybrid Networks), FAug (Feature Augmentation/Feature Learning), PL (Prompt-Learning), CRD (Cross Domain), ViT+CNN (Hybrid Networks), SDKD (Source Domain Knowledge Decoding), KD (Knowledge Distillation), SFDA (Source-Free DA), MML (Multi Modal Learning), and TTA (Test-Time Adaptation).
@@
Table 5: A summary of key features, advantages and challenges of different Transformers based methods in domain adaptation and domain generalization approaches.
Approaches Category Key Features Studies Advantages Challenges
Domain Adaptation Feature-Level adaptation Aligning feature distributions between domains DOT [90], TRANS-DA [93], DoT [94], SUDA [95], SAMB [96], DePT [97], CTTA [98] Reduces domain discrepancy May require complex feature engineering
Instance-Level adaptation Selecting/weighting data points relevant to target SF-OSDA [99] Enhances relevant feature learning Computationally intensive
Model-Level adaptation Modifying model architecture for better adaptability GE-ViT [39], TransDA[100], BeiT [101], TFC [102] Directly addresses model limitations Necessitates redesigning of model architecture
Hybrid Approaches Integrating feature, instance, and model adaptation techniques CDTRANS [87], TVT [88], SSRT [89], BCAT [103], PMTrans [91], UniAM [118], CoNMix [105], WinTR [106] Leverages strengths of multiple approaches Complexity in integrating different methods
Domain Generalization Multi-Domain Learning Training on multiple source Domains and capturing features invariant across multiple domains INDIGO [107] Improves feature robustness by leveraging diverse data Requires diverse source domains
Meta-Learning Meta-training and meta-testing to adapt to new domains with minimal data DPL [108], MoE [109], CSVPT [110] Enhances adaptation speed and effectiveness in new environments May require extensive meta-training
Regularization Imposing constraints during training and preventing overfit to source domains SDViT [111], CADG [112], RRLD [113], T3A [114] Boosts model robustness and ensures stable performance across domains Balancing regularization strength, implementation complexity, performance trade-offs
Data Augmentation Generates synthetic variations to simulate target domain characteristics TFS-ViT [115], ConvTran [116], AugDA[44], DAFormer[47] Provides diverse training scenarios, improving model adaptability Ensuring augmented data relevance
@@ 4.2 Action Recognition
Lin et al. [142] introduce ViTTA, a method enhancing action recognition models during test time without retraining. This approach focuses on feature distribution alignment, dynamically adjusting to match test set statistics with those of the training set. A key aspect is its applicability to both convolutional and transformer-based networks. ViTTA also enforces consistency in predictions over temporally augmented video views, a strategy that significantly improves performance in scenarios with distribution shifts, showcasing its effectiveness over previous test-time adaptation techniques. Q. Yan and Y. Hu’s research [143] introduces a UDA method tailored for skeleton behavior recognition, addressing the challenge of aligning source and target datasets in domain adaptation. Their method employs a spatial-temporal transformer framework with three flows—source, target, and source-target facilitating effective domain alignment and handling variations in joint numbers and positions across datasets. Key to this approach is the use of subsequence encoding and an attention mechanism that emphasizes local joint relationships, thereby enhancing the representation of skeleton behavior. Comprehensive testing on various skeleton datasets shows the superiority of their Spatial-Temporal Transformer-based DA (STT-DA) method, underscoring its effectiveness in managing the complexities of domain adaptation in skeleton behavior recognition. The concept of applying transformers for skeleton-based action recognition in DA is viewed as a promising and potentially impactful direction in this field of study [144].
@@ 4.3 Face Analysis
Face anti-spoofing (FAS) is a crucial aspect of biometric security systems, addressing the challenge of distinguishing between genuine and fake facial representations [145, 146]. Recently WACV 2023 research [147]introduced a new approach for FAS using ViTs. The authors proposed the Domain-invariant ViT (DiVT) which employs two specific losses to enhance generalizability. The losses include a concentration loss for learning domain-invariant representations by aggregating features of real face data, and a separation loss to differentiate each type of attack across domains. The study highlights the effectiveness of transformers in capturing long-range dependencies and globally distributed cues, crucial for FAS tasks. It also addresses the large model size and computational resource issues commonly associated with transformer models by adopting a lightweight transformer model, MobileViT. The proposed approach differs from previous methods by focusing on the core characteristics of real faces’ features and unifying attack types across domains, leading to improved performance and efficiency in FAS applications. In addition, researchers in [145] explored FAS in surveillance contexts, where image quality varies widely. The paper introduces an Adversarial DG Network (ADGN), which classifies training data into sub-source domains based on image quality scores. It then employs adversarial learning for feature extraction and domain discrimination, achieving quality-invariant features. The approach also integrates transfer learning to mitigate limited training data issues. This innovative method proved effective in surveillance FAS, as evidenced by its performance in the 4th Face Anti-Spoofing Challenge at CVPR 2023.
@@ 4.4 Medical Imaging
Another challenging area is medical image segmentation, which struggles due to the inherent variability of medical images [153]. To tackle the limited availability of training datasets, data-efficient ViTs were proposed. However, indiscriminate dataset combinations can result in Negative Knowledge Transfer (NKT). Addressing this, the introduction of MDViT, a multi-domain ViT with domain adapters, marked a significant advancement. This approach allowed for the effective utilization of varied data resources while mitigating NKT, showcasing superior segmentation performance even with an increase in domain diversity [154]. The robustness against adversarial attacks is a non-negotiable aspect of deep medical diagnosis systems. A novel CNN-Transformer hybrid model was introduced to bolster this robustness and enhance generalization. This model augmented image shape information in high-level feature spaces, smoothing decision boundaries and thereby improving performance on standardized datasets like MedMNIST-2D [155]. Liu et al.’s introduction of the Convolutional Swin-Unet (CS-Unet) represents a notable advance in medical image semantic segmentation. By integrating the Convolutional Swin Transformer (CST) block into transformers, they effectively combined multi-head self-attention with convolutions, providing localized spatial context and inductive biases essential for delineating organ boundaries. This model’s efficiency and effectiveness, particularly in its capability to surpass existing transformer-based methods without extensive pre-training, underscore the importance of localized spatial modeling in medical imaging [156]. A significant stride in domain generalization was achieved with the introduction of BigAug, a deep stacked transformation approach. This method applies extensive data augmentation, simulating domain shifts across MRI and ultrasound imaging. BigAug’s application of nine transformations to each image, validated across diverse segmentation tasks and challenge datasets, set a new standard, outperforming traditional augmentation and domain adaptation techniques in unseen domains [157].
@@
In reviewing the recent advancements in the application of ViTs for DA and DG in medical imaging, it becomes evident that ViTs are not only versatile but also increasingly effective in addressing domain-specific challenges. The studies surveyed indicate a significant shift towards models that are more adaptable to varying clinical environments, a crucial aspect of real-world medical applications. However, there is an observable need for further refinement in these models to ensure even greater accuracy and robustness, particularly in scenarios involving scarce or highly variable data. The success of models like EPVT and MDViT in enhancing generalization capabilities across diverse environments suggests a promising direction toward more domain-agnostic approaches. Nevertheless, the balance between model complexity and interpretability remains a key area for future exploration. As the field moves forward, there’s potential for integrating more advanced self-supervised learning techniques and exploring hybrid models that combine the strengths of both CNNs and ViTs. This could lead to a new generation of medical imaging tools that are not only more efficient in handling domain shifts but also more accessible and reliable for clinicians in varied healthcare settings.
@@
To conclude, this chapter has showcased the extensive utility of ViTs beyond image recognition tasks, highlighting their significant impact in areas such as semantic segmentation, action recognition, face analysis, and medical imaging. Figure 6 illustrates the categorization of studies utilizing Vision Transformers for domain adaptation and domain generalization across various tasks beyond image recognition. Beyond these fields, ViTs have proven to be highly adaptable and effective in sectors like precision agriculture and autonomous driving [161, 162]. These researches highlight that the potential of ViTs extends far beyond the initially discussed applications [163]. Their adaptability to different environments and challenges showcases a growing research interest in diverse fields. Future research could explore even more innovative applications, leveraging ViTs’ unique ability to handle complex visual tasks and distribution shifts across different industries.
Figure 6: Comprehensive categorization of studies leveraging vision transformers for domain adaptation and generalization across diverse tasks, including semantic segmentation [134], action recognition [144], face analysis [145], and medical imaging [150], extending beyond the image recognition tasks.
5 Conclusion, Discussion and Future Research Directions
@@
This comprehensive review looks at how modern vision networks, ViTs, are used in DA and DG methods to handle distribution shifts. Reviewing the research, we’ve observed that the transformer’s self-attention mechanisms play a pivotal role in generalizing and adapting to new, distribution-shifted samples, as evidenced by reviewing the experiment results of research in the known shifted datasets including ImageNet-A, ImageNet-C, and Stylized ImageNet. With respect to distribution shifts’ handling methods, for DA methods, we’ve organized the research into four categories: feature-level, instance-level, model-level adaptations, and hybrid approaches. Additionally, we’ve introduced a new category for studies that combine different strategies with ViTs to tackle distribution shifts. These hybrid networks show the use of various strategies alongside ViTs, leading us to categorize the papers based on these combined approaches. A similar methodological framework was applied to DG, wherein papers are classified based on multi-domain learning, meta-learning, regularization techniques, and data augmentation strategies. Our review is the first to catalog the use of ViTs in DA and DG tasks comprehensively. The tables provided organize and present an overview of the extant literature, demonstrating the burgeoning interest and application of ViTs across various domains. We observed that while the field is rapidly expanding, the literature still shows a sparsity. This has led us to implement additional categorizations for the diverse strategies ViTs employ in DA and DG, enhancing the readability and analytical depth for researchers. These two distinct methods of categorization, by considering various approaches, provide deeper insights and a broader perspective. Finally, we extend our review to applications beyond image recognition, showcasing the versatility of these methods in various domains, and highlighting the potential of ViTs in a broad range of real world applications particularly in critical safety and decision-making scenarios beyond the image recognition tasks.
In our discussion and exploration of various research works, it becomes evident that we are at the nascent stage of developing these modern vision networks. Researchers are increasingly focusing on the characteristics of ViTs across diverse deep learning scenarios. The challenges we faced in compiling and categorizing sparse references have been significant, particularly due to the rapid adoption and development of ViTs. While many papers claim the superiority of ViTs over CNNs, a balanced perspective considering both architectures is essential, depending on their robustness and the features crucial for specific applications. This aspect was evident in our exploration of robustness, where different factors important in generalization and stability were considered, with ViTs sometimes outperforming CNNs in extracting certain features.
While ViTs face challenges stemming from various factors, they are on a trajectory of improvement, much like CNNs were in their early stages. There is a noticeable rise in publications dedicated to modern vision models, emphasizing the advancements being made. As highlighted earlier, the attention mechanism inherent in ViTs significantly enhances their proficiency in handling distribution shifts, marking an advantage over CNNs in this aspect.
@@
In line with the focus of this study, this section outlines potential research areas in the field of tackling distribution shifts. The recommendations take into account the properties of ViTs, as well as DA and DG approaches for managing distribution shifts, and how they can be effectively integrated. ViTs have demonstrated the capacity to surpass previous state-of-the-art methods in certain contexts. However, their superiority in some tasks is not consistently overwhelming, partly due to their reliance on manually designed architectures, which may impede their adaptability, as explored by [52].
Despite the comprehensive nature of this review, it is noteworthy to mention the following limitations with respect to our research. Our review primarily focuses on studies using well-known benchmark datasets, which may limit the generalizability of our findings to other datasets and real-world scenarios not covered in the review. The datasets reviewed may not fully capture the diversity of real-world distribution shifts, potentially overlooking scenarios where ViTs might struggle. Additionally, the review is based on the current state of research, which may evolve with new methodologies and findings. Some promising methods might not have been included due to publication timing. Furthermore, while theoretical and experimental results are extensively reviewed, there is limited practical validation of ViTs in real-world applications within this paper. Certain assumptions were made regarding the performance and applicability of ViTs based on existing literature, which might not hold true in all practical scenarios.
In conclusion, this review not only aggregates and analyzes the role of ViTs within DA and DG frameworks but also outlines potential areas for future research, aimed at creating more robust and versatile deep learning models. As the first survey of its kind, it marks a significant step in understanding and advancing the capabilities of modern vision networks in handling distribution shifts, pointing towards a promising future in this dynamic field. Looking forward, the field faces substantial challenges such as the extensive data requirements and computational intensity of ViTs, alongside a need for real-world, application-specific datasets to validate new DA and DG approaches. To address these limitations and further enhance the applicability of ViTs, future research should focus on several key areas, including expanding the diversity of datasets used, improving practical validation in real-world applications, and keeping pace with evolving methodologies. In the following subsections, we will detail the challenges and outline future research directions accordingly.